# GraphRCG: Self-conditioned Graph Generation via Bootstrapped Representations

### v1

@@ 4.5. Representation Quality
4.6. Effect of Self-Conditioned Modeling
Our generation framework is conditioned on representations obtained from the representation generator, which encapsulates the graph distribution in a parametrized manner. On the other hand, previous works (Vignac et al., 2022) have also explored various strategies to capture and model distributions for the guidance of graph generation. For example, DiGress leverages three types of structural features: the number of cycles of different sizes, the spectral features, and the molecular characteristics (the valency of each atom and the molecular weight of the molecule) for graph generation. However, these are manually-designed features, which contain less information and require domain knowledge. In this subsection, we compare several variants of our framework with these structural features to evaluate the efficacy of self-conditioned modeling. In particular, we consider using the following features (representations) to replace our self-conditioned modeling module: (1) structural features proposed in DiGress, (2) representations of true training samples, which limit the uniqueness of generation, (3) a mixture of true representations, and (4) distributions learned by the Gaussian Mixture Model (GMM). As the structural features involve molecular information, we conduct experiments on the QM9 dataset with explicit hydrogens, which is a more complex setting with larger graphs.
Table 6. The ablation study results regarding self-conditioned guidance on the dataset Ego.
Dataset Ego
GraphRCG-Gradient 0.0134 0.0955 0.0464 0.0151
GraphRCG w/o $\mathcal{L}_{AG}$ 0.0088 0.0647 0.0252 0.0092
GraphRCG-fixed 0.0053 0.0653 0.0310 0.0125
GraphRCG 0.0015 0.0448 0.0183 0.0042
4.7. Effect of Self-Conditioned Guidance
In this subsection, we investigate the effect of self-conditioned guidance in our framework. We first replace the entire guidance strategy with a gradient-based method, which leverages computed gradients as guidance. Therefore, this variant does not require the training part in self-conditioned guidance, which means the graph generator does not involve any representation during training. For the second variant, we remove the alignment loss $\mathcal{L}_{AG}$ described in Eq. (10). Without this loss, the graph generator is not well-aligned with the representations produced by the representation generator, thereby affecting the guidance performance. For the third variant, we directly use the fixed representation, i.e., the clean representation, to guide generation. In this case, the step-wise guidance strategy is removed, resulting in the infeasibility of progressive guidance. We provide further details regarding the implementation of these variants in Appendix C.3. From the results presented in Table 6, we could first observe that our framework outperforms all other variants in most evaluation metrics, demonstrating the effectiveness of our self-conditioned guidance strategy. Moreover, GraphRCG-Gradient and GraphRCG w/o $\mathcal{L}_{AG}$ exhibit significantly poorer performance, as evidenced by higher scores across all metrics. This deterioration, particularly in Clus. and Orbit values, underscores the difficulty in preserving graph distributions without representation guidance or alignment loss. GraphRCG-fixed shows improved performance relative to the other variants, demonstrating the benefits of representation guidance even with a fixed one. Nevertheless, the results also indicate that the step-wise self-conditioned guidance is more beneficial for progressively guiding the generation process.
4.8. Comparison with Other Methods
In this subsection, we discuss and compare our framework GraphRCG with other diffusion-based generative models for graph data. We provide the comparisons regarding technical details in Table 5. Typically, existing diffusing methods for graph generation rely on either continuous or discrete diffusion. As a classific example, GDSS (Jo et al., 2022) adopts Gaussian transition kernels to perform continuous diffusion with a score-matching strategy. DiGress, on the other hand, performs discrete diffusion while considering categorical features of nodes and edges. Nevertheless, although continuous representations could capture complex structural patterns in graph distributions, they are less effective in generating discrete graph data. In contrast, our framework is capable of leveraging continuous diffusion to guide discrete diffusion, thereby combining the strengths of both and allowing for the generation of a wider range of graph structures. Regarding the convergence type, which refers to the pure noise state, we also combine both continuous and discrete noise to facilitate the optimization of both our representation generator and graph generator. Comparing the conditional generation type, our framework is conditioned on representations, which preserve richer information of complex distributions of each dataset. As a result, the generation process could significantly benefit from the guidance of representations.
5. Conclusion
In this work, we investigate the importance of capturing and utilizing distributions for graph generation. Specifically, we propose a novel self-conditioned generation, that encompasses self-conditioned modeling and self-conditioned guidance. Instead of directly learning from graph distributions, we encode all graphs into representations to capture graph distributions, which could preserve richer information. Our self-conditioned guidance module further guides the generation process in each timestep with representations with different degrees of noise. We conduct extensive experiments to evaluate our framework, and the results demonstrate the efficacy of our framework in graph generation.
Appendix A Noise Model
(12) $\mathbf{Q}_{t}^{\mathbf{X}}=\alpha^{t}\mathbf{I}+\beta^{t}\mathbf{1}_{a}(\mathbf{p}^{\mathbf{X}})^{\top},\ \ \text{and}\ \ \mathbf{Q}_{t}^{\mathbf{E}}$ $=\alpha^{t}\mathbf{I}+\beta^{t}\mathbf{1}_{b}(\mathbf{p}^{\mathbf{E}})^{\top}.$
[…]
Appendix B Self-conditioned Modeling
Algorithm 2 The training process of self-conditioned modeling.
0:  Optimized graph encoder and representation generator.
1:  for $i=1,2,\dotsc,T_{tr}$ do
2:     Sample $G=(\mathbf{X},\mathbf{E})$ from $\mathcal{G}$;
4:     $\mathbf{h}_{0}\leftarrow h_{\eta}(\mathbf{X},\mathbf{E})$;
6:     $\widetilde{\mathbf{h}}_{t}=f_{\gamma}\left(\mathbf{h}_{t},t\right)$;
[…]
10:     Optimize $h_{\eta}$ with $\mathcal{L}_{AR}$ in Eq. (3);
11:  end for
Appendix C Experimental Settings
C.1. Training Settings
@@
All experiments in our evaluation part are conducted on an NVIDIA A6000 GPU with 48GB of memory. For the specific parameters in each dataset, we follow the settings in DiGress (Vignac et al., 2022) and SparseDiff (Qin et al., 2023). We utilize the Adam optimizer (Kingma and Ba, 2015) for optimization. We also utilize the Xavier initialization (Glorot and Bengio, 2010). For the representation generator, we implement it as an RDM (Representation Diffusion Model) used in RCG (Li et al., 2023). The representation dimension size is set as 256 for all datasets. The learning rate of the RDM is set as $10^{4}$, and the weight decay rate is set as 0.01. The number of residual blocks in the RDM is set as 12. The batch size is various across datasets, adjusted according to the GPU memory consumption in practice. The total number of timesteps $T$ is set as 1,000. For the graph transformer architecture used to implement our graph generator, we set the number of layers as 8, with the hidden dimension size as 256.
C.2. Required Packages
@@
C.3. Ablation Study for Self-conditioned Guidance
@@
(14) $G_{t-1}\sim p_{\theta}(G_{t-1}|G_{t})p_{\eta}({\mathbf{h}}_{t}|G_{t-1}),$
where $p_{\eta}({\mathbf{h}}_{t}|G_{t-1})\propto\exp(-\lambda\langle\nabla_{G_{t}}\|\mathbf{h}_{t}-h_{\eta}(G_{t})\|_{2}^{2},G^{t-1}\rangle)$. The form is based on the classifier-guided DDIM sampling strategy in ADM (Dhariwal and Nichol, 2021). As a result, the training of the denoising network does not involve representation, which reduces the complexity during optimization. However, as the optimizations of the representation generator and the graph generator are separate, the guidance effect of representations cannot be guaranteed during sampling.
@@
Appendix D Visualization Results
@@
Figure 6. The generated graphs from the SBM dataset.
@@
Figure 7. The generated graphs from the Planar dataset.

### v2

@@ C.1 Comparison with Other Methods
Appendix D Experimental Settings
Table 6: Detaile statistics of generic datasets used in our experiments.
Name Graph number Node range Edge range
Planar 200 [64, 64] [346, 362]
SBM 200 [44, 187] [258, 2258]
Ego 757 [50, 399] [112, 2124]
D.1 Dataset Details
For generic datasets, we consider SBM, Planar, and Ego. In particular, the two synthetic datasets, SBM and Planar, are obtained following the settings in SPECTRE [33]. The Ego dataset setting follows SparseDiff [36].
  • The SBM dataset contains 2022 Stochastic Block Model graphs. Each graph has 2 to 5 communities, each of which has 20 to 40 nodes. The inter-community edge probability is 0.3 and the intra-community edge probability is 0.05.
  • The Planar dataset contains 200 planar graphs, each with 64 nodes. The graphs are generated via Delaunay triangulation on a set of randomly placed points.
  • The Ego dataset contains 757 graphs, each with 50 399 nodes, sampled from the Citeseer Network Dataset [11].
The detailed statistics of these datasets are provided in Table 6.
For molecular datasets, we consider ZINC250k and QM9 following GraphARM [24]. Particularly, QM9 consists of molecular graphs with a maximum of nine heavy atoms that could be treated with implicit or explicit hydrogens. We provide the details of dataset statistics in Table 7. ZINC-250k consists of approximately 250,000 drug-like molecules, each of which has up to 38 atoms. The dataset contains nine atom types and three edge types.
Table 7: Detaile statistics of molecular datasets used in our experiments.
Dataset Number of graphs Node range Number of node types Number of edge types
QM9 133,885 [1, 9] 4 3
ZINC250k 249,455 [6, 38] 9 3
D.2 Training Settings
@@
All experiments in our evaluation part are conducted on an NVIDIA A6000 GPU with 48GB of memory. For the specific parameters in each dataset, we follow the settings in DiGress [52] and SparseDiff [36]. We utilize the Adam optimizer [23] for optimization. We also utilize the Xavier initialization [12]. For the representation generator, we implement it as an RDM (Representation Diffusion Model) used in RCG [27]. The representation dimension size is set as 256 for all datasets. The learning rate of the RDM is set as $10^{4}$, and the weight decay rate is set as 0.01. The number of residual blocks in the RDM is set as 12. The batch size is various across datasets, adjusted according to the GPU memory consumption in practice. The total number of timesteps $T$ is set as 1,000. For the graph transformer architecture used to implement our graph generator, we set the number of layers as 8, with the hidden dimension size as 256. We provide our code in the supplementary materials.
D.3 Evaluation Metrics
For molecular graph generation, following previous works [21, 24], we evaluate the generated molecular graphs with several key metrics: (1) Frechet ChemNet Distance (FCD) [35], which quantifies the discrepancy between the distributions of training and generated graphs, based on the activation values obtained from the penultimate layer in ChemNet. (2) Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) MMD [5], which measures the Maximum Mean Discrepancy (MMD) between the generated and the test molecules, considering both node and edge features. (3) Validity, which refers to the proportion of generated molecules that are structurally valid without requiring valency adjustments. (4) Uniqueness, which measures the proportion of unique molecules that are distinct from each other.
D.4 Required Packages
@@
D.5 Ablation Study for Self-conditioned Guidance
@@
$G_{t-1}\sim p_{\theta}(G_{t-1}|G_{t})p_{\eta}({\mathbf{h}}_{t}|G_{t-1}),$ (15)
where $p_{\eta}({\mathbf{h}}_{t}|G_{t-1})\propto\exp(-\lambda\langle\nabla_{G_{t}}\|\mathbf{h}_{t}-h_{\eta}(G_{t})\|_{2}^{2},G^{t-1}\rangle)$. The form is based on the classifier-guided DDIM sampling strategy in ADM [7]. As a result, the training of the denoising network does not involve representation, which reduces the complexity during optimization. However, as the optimizations of the representation generator and the graph generator are separate, the guidance effect of representations cannot be guaranteed during sampling.
@@
Table 8: The performance of various methods on three generic datasets.
Model Community Cora Enzymes
Deg. Clus. Orbit Deg. Clus. Orbit Deg. Clus. Orbit
SPECTRE 0.048 0.049 0.016 0.021 0.080 0.007 0.136 0.195 0.125
GDSS 0.045 0.086 0.007 0.160 0.376 0.187 0.026 0.061 0.009
DiGress 0.047 0.041 0.026 0.044 0.042 0.223 0.004 0.083 0.002
GraphARM 0.034 0.082 0.004 0.273 0.138 0.105 0.029 0.054 0.015
GraphRCG 0.040 0.053 0.029 0.038 0.036 0.173 0.004 0.079 0.002
Appendix E Additional Results
In this section, we consider additional experiments on three generic datasets. As our framework is based on discrete graph diffusion, we mainly follow the dataset settings used in DiGress [52] and the following work SparseDiff [36] for the three generic datasets. We consider the following datasets: Community [57], Cora [42], and Enzymes [41], order to further improve the integrity of our evaluation. We provide the results in Table 8. The results of other baselines are obtained from GraphARM [24]. From the results, we could observe that our framework also achieves competitive results, compared to other methods on various datasets. The performance is particularly better on dataset Cora and Enzymes, indicating that our framework is generalizable to various datasets.
Appendix F Visualization Results
In this section, we provide additional visualization results from the experiments of generic graph generation. Particularly, we sample different generation results from the SBM and Planar datasets. The results are provided in Fig. 5 and Fig. 6.
@@ Appendix F Visualization Results
Figure 5: The generated graphs from the SBM dataset.
@@
Figure 6: The generated graphs from the Planar dataset.
(a) SBM
(b) Planar
Figure 7: The t-SNE visualization of the Planar and SBM datasets, along with generated representations by our framework. Each point denotes the representation of a training graph sample or from our representation generator.
Appendix G Representation Quality
In this section, we provide visualization results to evaluate the quality of the generated representations. We present the t-SNE plot of representations from training samples and our representation generator on datasets SBM and Planar in Fig. 7. From the visualization results, we could observe that our representation generator could faithfully capture the graph distributions from training samples. In other words, the generated representations closely align with training graph distributions. More importantly, our framework can generate representations that slightly deviate from the training samples in specific directions. This indicates that our framework is capable of discovering novel graph distributions that are not present in training samples, while still ensuring the validity of these representations for further guidance during generations.
Appendix H Limitation
In this section, we discuss the potential limitations presented in our framework. Although our framework is capable of generating graphs conditioned on the representations provided by our representation generator, it also means that the optimization quality of the representation generator is critical for graph generation. In other words, if the representation generator is not well-trained, the output representations could be detrimental to graph generation. In addition, the learned representations are not evaluated for their quality, as there are not ground truths for evaluation. A feasible solution is to create a dataset that is directly generated from representations. In this manner, the representations learned in our framework could be evaluated for their quality and similarity to the ground truths.
Appendix I Broader Impact
As our datasets only involve datasets, the information in these datasets is ensured to be anonymized. We only use datasets from public releases, and thus we infer that this work does not have negative social impacts.